Calculate the principal sum which, if deposited at 9.5% compound interest, will grow to £400 after three years.



What is asked:



What are given:



What is the answer:

1
Expert's answer
2021-11-15T14:30:13-0500

Here,

What is asked: principal amount (P)

What are given:

  • Interest rate (r)= 9.5% compounded
  • Future value (F)= £400
  • Timer period (t)= 3 years


According to the given information, answer (principal amount) is:

"=\frac{F}{(1+r)^t}\\=\frac{400}{(1+0.095)^3}\\\frac{400}{1.31293228}\\=304.661539\\=305"